OpenStudy (anonymous):

Which conic section does this equation represent? -4x^2-9x+4y^2+8x=16 A. Circle B. Ellipse C. Parabola D. Hyperbola

OpenStudy (anonymous):

-4x^2-9x+8x+4y^2=16 -4x^2-x+4y^2=16 -x^2-1x/4+y^2=4 -{x^2+x/4+(1/8)^2-(1/8)^2}+y^2=4 -(x+1/8)^2+1/64+y^2=4 -(x+1/8)^2+y^2=4-1/64=255/64 y^2/(255/64)-(x+1/8)^2/(255/64)=1 which is of the form y^2/a^2-x^2/b^2=1 which is a hyperbola.